Use the value of the discriminant to determine the number and type of roots for the equation: 2x^2-7x+9=0.
Thanks

Answer by venugopalramana(3286)

Use the value of the discriminant to determine the number and type of roots for the equation: 2x^2-7x+9=0.
DISCRIMINANT =B^2-4AC=7^2-4*2*9=49-72= - VE...SO ROOTS ARE IMAGINARY